OpenStudy (anonymous):

The length of a rectangle is 11 m and its width is 8 m. what is the area of ½ of the rectangle.

OpenStudy (anonymous):

What's the area of the full rectangle?

OpenStudy (anonymous):

\( Length\space \times \space Width = Area\) \( \frac{Length\space \times \space Width}{2} = Answer\)